# Reassessing Sub-Neptune Structure, Radii, and Thermal Evolution

### v1

@@ 2.1 Physical Structure
where $r$, $\rho$ and $P$ represent radius, density and pressure at a certain mass shell, respectively, and $m$ denotes the mass enclosed within that mass shell. To improve boundary condition assessments and enhance computational efficiency, we adopt a higher spatial resolution by using smaller mass shells ($dm$) at larger radii. The number of cells in each structural layer is carefully selected to ensure that the physical structures converge across different resolution choices. The governing equations are solved using a relaxation method, where the initial guess is taken from the converged solution of the preceding timestep.
@@
Figure 1: This schematic is designed to emphasize the essential interfaces, terminologies, and structural components. The relative size of each layer is not to scale. Key locations include the radiative-convective boundary (RCB), the mantle-envelope boundary (MEB), the core-mantle boundary (CMB) and the phase change interfaces (PCI) within both the mantle and core. Additionally, the diagram marks the olivine-to-magnesium perovskite transition at 25 GPa, the optical transit radius at 20 mbar, and the thermosphere, characterized by significant EUV absorption. Our definitions of the rock/iron core, planetary interior, and atmosphere are also provided for clarity, as these may differ from those used in other studies.
@@ 2.3 Radiative Atmosphere
Sub-Neptunes’ intrinsic flux is usually negligible compared to stellar bolometric flux. Compared to giant planets, sub-Neptunes have cooler interiors due to their smaller energy reservoirs and relatively fast cooling rates, which are facilitated by their lower gravity. This combination of low RCB temperature and rapid cooling leads to a deeper RCB location (at a higher pressure), making the location of the RCB important in the radius calculation. Our numerical model estimates the RCB location by finding the joint point in pressure space between the adiabatic temperature profile of the H/He envelope and the radiative atmospheric profile. The RCB pressure of a sub-Neptune undergoes significant changes over time, ranging from around $\sim$ 10 bar in its early stages to 10 kbar in ages exceeding a few Gyrs. As a sub-Neptune ages, its internal temperature decreases, resulting in a reduction of its intrinsic luminosity. Consequently, the disparity between the intense flux emitted by the host star and the diminishing intrinsic flux of the planet widens, causing the RCB to extend deeper into the planet. In Figure 2, we show the locations of the RCBs on the T-P profiles of a sub-Neptune in cross shapes (bottom), and the thermal contraction of the interior in solid curves (top left).
@@
The temperature profile is primarily influenced by the opacity ratio $\gamma=\kappa_{\nu}/\kappa_{\rm{th}}$ where $\kappa_{\nu}$ and $\kappa_{\rm{th}}$ are the opacities to the visible stellar radiation and outgoing thermal radiation, respectively. We assume $\kappa_{th}$ to be a constant value of 0.02 $\rm{cm^{2}\ g^{-1}}$. To determine $\kappa_{\nu}$, we compare the analytical model’s $T-P$ profile to full radiative-convective equilibrium models for solar composition atmospheres (e.g., Marley & McKay, 1999; Fortney et al., 2005, 2020; Marley & Robinson, 2015; Thorngren et al., 2019; Ohno & Fortney, 2023), fitting the best values as a function of the incident bolometric flux. The values show weak dependence on surface gravity and intrinsic luminosity from the interior, as the radiative transfer is primarily driven by re-emitted stellar energy. Note that these $\kappa$ terms do not enter into the thermal evolution calculation, just into assessing shape of the radiative $T-P$ profile.
The temperature profiles of the radiative atmosphere typically exhibits two distinct isothermal regions: an outer “skin temperature” layer, which is optically thin to both incoming optical radiation and outgoing infrared radiation, and a deeper layer, characterized by the absorption of incoming optical radiation that is in equilibrium with the optically thick infrared radiation. This is shown in Figure 2 (dashed curves in bottom left panel).
@@
Figure 2: An example of the evolutionary trajectory for a 4 $M_{\oplus}$ planet with 2% H/He mass irradiated with 10 $F_{\oplus}$. The top left panel shows the planetary radii at the RCB (solid), 20 mbar (dashed), and 1 nbar (dotted) pressure levels. The top right panel presents the energy budget as the luminosities at each boundary layer, including the mantle’s phase change interface (PCI), and the contribution from the radiogenic heat. The bottom panels display the temperature-pressure profile at various ages. In the bottom panels, we mark the radiative-convective boundary (RCB, cross), mantle-envelope boundary (MEB, triangle), and core-mantle boundary (CMB, circle). In the bottom left panel, the radiative atmosphere is shown in dashed colors and the isothermal rock/iron core assumed in earlier models is represented with a dotted line. The bottom right panel zooms in on the rock/iron interior, focusing on its detailed thermal and structural evolution. It highlights phase transitions within the planet’s core and mantle. For comparison, we include the mantle solidus (dashed black), liquidus (dotted black), and iron melting temperature (dotted green).
@@ 2.4 Metals in the Envelope and Atmosphere
Beyond a setup for pure H/He envelopes in the previous section, we also investigate the role of metallicity on planetary evolution and structure.
First, we utilize an atmosphere grid for 50 times solar metallicity to assess the intrinsic luminosity (Lopez & Fortney, 2014) at the metal-rich conditions that may be appropriate from some sub-Neptunes (Fortney et al., 2013; Benneke et al., 2024). The format is similar to that of our default solar metallicity grid. Moreover, to account for the metals in the envelope, we choose to mix in a $Z$ mass fraction of water in the envelope. We use the additive volume law to obtain the thermodynamic properties, including the specific entropy and density at a certain pressure and temperature. A correction term is incorporated in calculating the entropy of the mixture to recover the ideal mixing case at low pressures and temperatures (Chabrier et al., 2019).
We use the water equation of state of Mazevet et al. (2019), which covers a wide range of physical states, including precisely calculated liquid, gas, and plasma phases and approximated ice and superionic phases. At 50$\times$ solar, the hydrogen, helium and water mass fractions are X = 0.4317, Y = 0.1638 and Z = 0.4045, respectively. (The hydrogen to helium mass ratio is the same as our pure H/He mixture.)
These parameters yield an increase in the mean molecular weight of $\mu=3.58$ (from 2.35) for the molecule-dominated region including the envelope and lower radiative atmosphere and $\mu=1.93$ (from 1.28) for the upper atmosphere where $\rm{H_{2}O}$ is dissociated into atomic H and OH radicals. To account for the non-ideal effect in the deep atmosphere and photodissociation in the upper atmosphere, we apply the same approaches as used in the pure H/He case. Unless specified, our numerical results are calculated with our default pure H/He setup.
@@ 2.5 Initial States from a Boil-Off Calculation
To compute the initial contraction timescale, we assume thermal equilibrium between the envelope and mantle, thus neglecting energy contributions from the rock/iron core and setting $L_{\rm{MEB}}=0$ in Eq. 15. This assumption holds for most model planets, since the rock/iron core contributes minimally to the total energy with the gravitational binding energy of the envelope dominating. However, for planets with extremely low envelope masses ($\sim$ 0.1%), the envelope’s energy reservoir $E_{\rm{env}}$ can be much smaller than that of the rock/iron core $E_{\rm{ric}}$, making the contraction timescale sensitive to core and mantle heat production. This includes radiogenic heating from ${}^{26}\mathrm{Al}$ and other heating sources not included in the model (which are likely on the order of $E_{\rm{ric}}$), such as solid mass accretion and rock/iron differentiation, which are beyond the scope of this model. For these cases, our above assumption likely underestimates the interior temperature, leading to partial core solidification at $t=0$. To correct for this, we assume a higher initial envelope specific entropy, ensuring a fully molten interior state at the beginning of thermal evolution.
@@ 2.6 Iron Core
where the reference temperature $T_{\rm{ref}}$ and the coefficients $a$ and $b$ are summarized in Table 2.7. To incorporate the effects of impurities, primarily sulfur, we adjust the melting temperature of pure iron by a specific fraction. Previous studies (Zhang & Rogers, 2022) on Earth’s interior evolution suggest using 80% of the pure iron melting temperature to accurately match the size of Earth’s inner liquid iron, which we adopt as our default value for super-Earths and vary to examine its influence on the solidification timescale in Section 3.4.
Given the interdependence of the physical states and temperature profile, we employ an iterative relaxation method for self-consistent calculations, initializing the physical states with the converged solution from the previous timestep. This iterative approach is crucial due to the sensitive nature of the solidification process.
Density at each mass shell in each compositional layer is derived from the EOS, which relates density to pressure. By integrating Eq. 1 and 2, we obtain the radius and pressure distributions. For computational efficiency, we have neglected the temperature-dependent thermal pressure in calculating the rock/iron density when generating the model results. However, in the full evolution model, we implemented the thermal pressure terms in the EOS using the Mie-Grüneisen-Debye model, with the parameters sourced from the references listed in Table 2.7.
This simplification leads to an overestimation of density and an underestimation of temperature at low pressure levels and high temperatures by less than 5%. The resulting impact on the planetary radius is minimal, typically on the order of 0.1%, which is insignificant since the overall radius of most sub-Neptunes is primarily determined by the H/He envelope and radiative atmosphere, rather than the rock/iron core.
@@
where $T_{\rm{CMB,c}}$, $\alpha_{\rm{CMB,c}}$, $\rho_{\rm{CMB,c}}$ are the temperature, thermal expansivity and density of the iron core at the CMB. By tracking $T_{\rm{CMB,c}}$, we can accurately evolve the iron core’s thermal state along the adiabat. The second term in the middle, associated with the pressure change at the CMB, $\dot{P}_{\rm{CMB}}$, is updated during each hydrostatic calculation. Although this pressure change typically evolves over a long timescale and contributes minimally to the energy budget during the thermal contraction phase, it becomes significantly impactful during periods of boil-off. Rapid depressurization due to the removal of envelope mass can substantially influence the interior energy budget, highlighting the necessity of this decompression term.
@@ 2.7 Silicate Mantle
For the solid mantle, the dominant mineral types vary with pressure. Below 25 GPa, the lower mantle consists primarily of Mg-perovskite ($\rm{MgSiO_{3}}$), while the upper solid mantle is composed of olivine ($\rm{Mg_{2}SiO_{4}}$). The EOS for Mg-perovskite is based on Keane’s formulation (Keane, 1954), which reduces uncertainties at high pressures compared to the BM3 model:
@@
where $K_{0}$, $K^{\prime}_{0}$, $K_{0}^{\prime\prime}$ and $\rho_{0}$ are the model parameters.
@@
The solidus for the upper mantle (pressures $<$ 10 GPa) is based on experiments on peridotite (Hirschmann, 2000). Since the liquidus data from the same experiments carry significant uncertainty, we assume the liquidus is 200 K above the solidus. At higher pressures, the solidus and liquidus are derived from experimental data on chondrites (Andrault et al., 2011). For pressures exceeding 300 GPa, where experimental data is unavailable, the liquidus is taken from the ab initio melting curve of pure Mg-perovskite (Belonoshko et al., 2005), with the solidus assumed to be 85% of the liquidus temperature. The T-P relationship of the solidus and liquidus are shown in the bottom right panel of 2. We integrate the local melt fraction $\phi$ over each mass shell to determine the total masses of the solid and molten silicates, denoted as $M_{s}$ and $M_{m}$, respectively.
The temperature difference across the interface between the two solid layers (the upper and lower solid mantle) is assumed to be zero, meaning they evolve together as a single thermodynamic entity. We evolve the magma ocean and solid mantle separately due to their contrasting rates of heat transport. The temperature jump at the PCI governs the energy transfer rate between the solid and liquid layers.
@@
where $M_{\rm{Si}}=M_{s}+M_{m}$ represents the total silicate mass. In these equations, the rate of change in the potential temperatures of the magma ocean, $T_{m}$, and solid mantle, $T_{s}$, are evaluated similarly to Eq. 25. $T_{c}$, $T_{m}$ and $T_{s}$ are evolved simultaneously with $s$ and $M_{\rm{env}}$ using the ODE solver, based on Eq. 24, 29 and 30. On the right hand sides, the energy transport rates ($L_{\rm{CMB}}$, $L_{\rm{PCI}}$ and $L_{\rm{MEB}}$) at each boundary layers are included, alongside heat production from silicate solidification $L_{\rm{Si}}=l_{\rm{Si}}\dot{M}_{\rm{s}}$ and radioactive decay $L_{\rm{radio}}$. An example of the evolutionary trajectory of these luminosities is shown in the top right panel of Figure 2. $L_{\rm{radio}}$ is evaluated with:
@@
The heat transport process in the mantle is dominated by mantle convection. To evaluate the luminosity at each boundary layer, we employ parameterized convection, which models the process using scaling laws derived from fluid dynamics instead of solving the full set of fluid dynamic equations. This approach has been widely used for Earth and other solar system bodies (e.g., McKenzie & Richter, 1981; Nimmo & Stevenson, 2000; Nimmo et al., 2004).
@@
where $k$ is thermal conductivity of the mantle, whose value varies depending on the physical state and the pressure level. $\Delta T_{\rm{CMB}}=T_{\rm{CMB},c}-T_{\rm{CMB},m}$ sets the temperature gradient at the boundary. The thickness of the thermal boundary layer $\delta$ is parameterized using scaling laws. Two regimes are considered here. The mantle convects in the so-called “hard turbulence” regime (Shraiman & Siggia, 1990) in the liquid state ($\phi>0.5$), because of the low viscosity and large thickness of the magma ocean. For the solid convection ($\phi\leq 0.5$), $\delta$ is evaluated using the scaling law for the local Rayleigh number $Ra$, which approaches the critical Rayleigh number $Ra_{c}$ at the thermal boundary (Nimmo & Stevenson, 2000; Nimmo et al., 2004). These give:
$\delta=\begin{dcases}d/Nu=d/(0.3Ra^{2/7}),&\text{if }\phi>0.5\\ \left[\frac{Ra_{c}\kappa_{m}\eta}{\rho g\alpha\Delta T_{\rm{CMB}}}\right]^{1/3},&\text{if }\phi\leq 0.5\\ \end{dcases}$ (33)
@@
where $\rho$ is local density, $g$ is local gravity, $\alpha$ is local thermal expansivity, $\kappa_{m}$ is the thermal diffusivity of the mantle, $\eta$ is the dynamic viscosity of the mantle, and $d$ is the thickness of the magma ocean. The Nusselt number $Nu$ parameterizes the increased heat transfer due to convection and depends on $Ra$.
@@
The parameters $\eta_{0}$, $E_{a}$, $V_{a0}$ and $T_{0}$ are provided in Table 2.7. The transition from the low-viscosity magma (with $100\ \rm{Pa\ s}$) to the highly viscous solid phase (with $10^{22}\ \rm{Pa\ s}$) is quite abrupt. To capture this behavior, we adopt a hyperbolic switch function $w$ with a transition width $\Delta\phi=0.1$ and a critical melt fraction $\phi_{\rm{crit}}=0.5$ (Zhang & Rogers, 2022):

### v2

@@ 2.1 Physical Structure
where $r$, $\rho$ and $P$ represent radius, density and pressure at a certain mass shell, respectively, and $m$ denotes the mass enclosed within that mass shell. To improve boundary condition assessments and enhance computational efficiency, we adopt a higher spatial resolution by using smaller mass shells ($dm$) at larger radii. The number of cells in each structural layer is carefully selected to ensure that the physical structures converge across different resolution choices. The governing equations are solved using a relaxation method, where the initial guess is taken from the converged solution of the preceding timestep. The relaxation method follows an iterative approach, repeatedly integrating Eq. 1 and Eq. 2 until the relative errors are reduced to below 0.1%.
@@
Figure 1: This schematic is designed to emphasize the essential interfaces, terminologies, and structural components. The relative size of each layer is not to scale. Key locations include the radiative-convective boundary (RCB), the mantle-envelope boundary (MEB), the core-mantle boundary (CMB) and the phase change interfaces (PCI) within both the mantle and core. Additionally, the diagram marks the olivine-to-magnesium perovskite transition at 250 kbar (25 GPa), the optical transit radius at 20 mbar, and the thermosphere, characterized by significant EUV absorption. Our definitions of the rock/iron core, planetary interior, and atmosphere are also provided for clarity, as these may differ from those used in other studies.
@@ 2.3 Radiative Atmosphere
Sub-Neptunes’ intrinsic flux is usually negligible compared to stellar bolometric flux. Compared to giant planets, sub-Neptunes have cooler interiors due to their smaller energy reservoirs and relatively fast cooling rates, which are facilitated by their lower gravity. This combination of low RCB temperature and rapid cooling leads to a deeper RCB location (at a higher pressure), making the location of the RCB important in the radius calculation. Our numerical model estimates the RCB location by finding the joint point in pressure space between the adiabatic temperature profile of the H/He envelope and the radiative atmospheric profile.
This approach results in a deeper RCB location compared to models that assume an isothermal atmosphere. While this depth still remains a slight underestimate relative to the true RCB determined by radiative-convective equilibrium modeling (Ohno & Fortney, 2023), the agreement is quite good. An example is shown in Figure 2, where we compare the T-P profiles and RCB locations from our model (red) with those from radiative-convective modeling (black) and an isothermal atmosphere model (gray dashed).
The RCB pressure of a sub-Neptune undergoes significant changes over time, ranging from around $\sim$ 10 bar in its early stages to 10 kbar in ages exceeding a few Gyrs. As a sub-Neptune ages, its internal temperature decreases, resulting in a reduction of its intrinsic luminosity. Consequently, the disparity between the intense flux emitted by the host star and the diminishing intrinsic flux of the planet widens, causing the RCB to extend deeper into the planet. In Figure 3, we show the locations of the RCBs on the T-P profiles of a sub-Neptune in cross shapes (bottom), and the thermal contraction of the interior in solid curves (top left).
@@
The temperature profile is primarily influenced by the opacity ratio $\gamma=\kappa_{\nu}/\kappa_{\rm{th}}$ where $\kappa_{\nu}$ and $\kappa_{\rm{th}}$ are the opacities to the visible stellar radiation and outgoing thermal radiation, respectively. We assume $\kappa_{th}$ to be a constant value of 0.02 $\rm{cm^{2}\ g^{-1}}$. To determine $\kappa_{\nu}$, we compare the analytical model’s $T-P$ profile to full radiative-convective equilibrium models for solar composition atmospheres (e.g., Marley & McKay, 1999; Fortney et al., 2005, 2020; Marley & Robinson, 2015; Thorngren et al., 2019; Ohno & Fortney, 2023), fitting the best values as a function of the incident bolometric flux. The values show weak dependence on surface gravity and intrinsic luminosity from the interior, as the radiative transfer is primarily driven by re-emitted stellar energy. Note that these $\kappa$ terms do not enter into the thermal evolution calculation, just into assessing shape of the radiative $T-P$ profile. The fit demonstrates very good agreement between the analytical model (red) and the numerical model (black) in Figure 2.
The temperature profiles of the radiative atmosphere typically exhibits two distinct isothermal regions: an outer “skin temperature” layer ($T_{\rm{skin}}$), which is optically thin to both incoming optical radiation and outgoing infrared radiation, and a deeper layer ($T_{\rm{deep}}$), characterized by the absorption of incoming optical radiation that is in equilibrium with the optically thick infrared radiation. This is shown in Figure 2.
@@
Figure 2: A close-up view of the top of the convective H/He envelope and radiative atmosphere. For a complete view, see the bottom left panel of Figure 3. We compare the $T-P$ profiles and RCB locations derived from non-gray 1D radiative-convective atmosphere modeling (black) with those from our implementation (red). For reference, we also include the isothermal temperature profile (gray dashed) commonly used in previous sub-Neptune models. This plot is for a planet with $T_{\rm{int}}=150$K, $F_{\rm{bol}}=100F_{\oplus}$ and gravity of 200 $cm\ s^{-2}$. See text for discussion.
Figure 3: An example of the evolutionary trajectory for a 4 $M_{\oplus}$ planet with 2% H/He mass irradiated with 10 $F_{\oplus}$. The top left panel shows the planetary radii at the RCB (solid), 20 mbar (dashed), and 1 nbar (dotted) pressure levels. The top right panel presents the energy budget as the luminosities at each boundary layer, including the mantle’s phase change interface (PCI), and the contribution from the radiogenic heat. Planetary age is defined in the model as the time elapsed since the onset of inner disk dispersal. The bottom panels display the temperature-pressure profile at various ages. In the bottom panels, we mark the radiative-convective boundary (RCB, cross), mantle-envelope boundary (MEB, triangle), and core-mantle boundary (CMB, circle). In the bottom left panel, the radiative atmosphere is shown in dashed colors and the isothermal rock/iron core assumed in earlier models is represented with a dotted line. The bottom right panel zooms in on the rock/iron interior, focusing on its detailed thermal and structural evolution. It highlights phase transitions within the planet’s core and mantle. For comparison, we include the mantle solidus (dashed black), liquidus (dotted black), and iron melting temperature (dotted green).
@@ 2.4 Metals in the Envelope and Atmosphere
Beyond a setup for pure H/He envelopes in the previous section, we also investigate the role of metallicity on planetary evolution and structure. Our model is capable of simulating metal-enhanced evolution for up to 100 times solar metallicity. In this study, we default our metal-enhanced models to 50 times solar metallicity.
First, we expand our atmosphere grids to assess the intrinsic luminosity (Lopez & Fortney, 2014; Ohno & Fortney, 2023) at the metal-rich conditions that may be appropriate from some sub-Neptunes (Fortney et al., 2013; Benneke et al., 2024). The format is similar to that of our default solar metallicity grid. We interpolate within these atmosphere grids, covering [1, 3, 10, 30, 50, 100] times solar metallicity, to evaluate the luminosity at various metallicity values. Moreover, to account for the metals in the envelope, we choose to mix in a $Z$ mass fraction of water in the envelope. We use the additive volume law to obtain the thermodynamic properties, including the specific entropy and density at a certain pressure and temperature. A correction term is incorporated in calculating the entropy of the mixture to recover the ideal mixing case at low pressures and temperatures (Chabrier et al., 2019).
We use the water equation of state of Mazevet et al. (2019), which covers a wide range of physical states, including precisely calculated liquid, gas, and plasma phases and approximated ice and superionic phases.
The mass fractions and mean molecular weights at a given metallicity are self-consistently calculated. At 50$\times$ solar, the hydrogen, helium and water mass fractions are X = 0.4317, Y = 0.1638 and Z = 0.4045, respectively. (The hydrogen to helium mass ratio is the same as our pure H/He mixture.) These parameters yield an increase in the mean molecular weight of $\mu=3.58$ (from 2.35) for the molecule-dominated region including the envelope and lower radiative atmosphere and $\mu=1.93$ (from 1.28) for the upper atmosphere where $\rm{H_{2}O}$ is dissociated into atomic H and OH radicals. To account for the non-ideal effect in the deep atmosphere and photodissociation in the upper atmosphere, we apply the same approaches as used in the pure H/He case. Unless specified, our numerical results are calculated with our default pure H/He setup.
@@ 2.5 Initial States from a Boil-Off Calculation
To compute the initial contraction timescale, we assume thermal equilibrium between the envelope and mantle, thus neglecting energy contributions from the rock/iron core and setting $L_{\rm{MEB}}=0$ in Eq. 15. This assumption holds for most model planets, since the rock/iron core contributes minimally to the total energy with the gravitational binding energy of the envelope dominating. However, for planets with extremely low envelope masses ($\sim$ 0.1%), the envelope’s energy reservoir $E_{\rm{env}}$ can be much smaller than that of the rock/iron core $E_{\rm{ric}}$, making the contraction timescale sensitive to core and mantle heat production. This includes radiogenic heating from ${}^{26}\mathrm{Al}$ and other heating sources not included in the model (which are likely on the order of $E_{\rm{ric}}$), such as solid mass accretion and rock/iron differentiation, which are beyond the scope of this model. For these cases, our above assumption likely underestimates the interior temperature, leading to partial core solidification at $t=0$. To correct for this, we assume a higher initial envelope specific entropy, ensuring a fully molten interior state at the beginning of thermal evolution. The potential role of core-mantle differentiation is further discussed in Section 4.3.
@@ 2.6 Iron Core
where the reference temperature $T_{\rm{ref}}$ and the coefficients $a$ and $b$ are summarized in Table 2.7. To incorporate the effects of impurities, primarily sulfur, we adjust the melting temperature of pure iron by a specific fraction. Previous studies (Stevenson et al., 1983; Stixrude, 2014; Zhang & Rogers, 2022) on Earth’s interior evolution suggest using 80% of the pure iron melting temperature to accurately match the size of Earth’s inner liquid iron, which we adopt as our default value for super-Earths and vary to examine its influence on the solidification timescale in Section 3.4.
Given the interdependence of the physical states, temperature profiles and physical structures, we employ the iterative relaxation method described in Section 2.1 for self-consistent calculations of all these variables. This relaxation approach is crucial due to the sensitive nature of the solidification process.
Density at each mass shell in each compositional layer is derived from the EOS, which relates density to pressure at a reference temperature $T_{0}$. The density calculation also depends weakly on the temperature profile due to thermal perturbations. Thus, the full EOS is expressed as:
$P(\rho,T)=P(\rho,T_{0})+P_{\rm{th}}(\rho,T)-P_{\rm{th}}(\rho,T_{0})$ (23)
where $P_{\rm{th}}$ represents the temperature-dependent thermal pressure. The thermal pressure terms are calculated using the Mie-Grüneisen-Debye model, expressed as:
$P_{\rm{th}}(\rho,T)=\gamma\rho E_{\rm{th}}(\rho,T)$ (24)
where the parameters for $E_{\rm{th}}$ are sourced from the references listed in Table 2.7. The density $\rho(P,T)$ is determined through linear interpolation of Eq. 23. By integrating Eq. 1 and 2, we obtain the radius and pressure distributions.
For computational efficiency, we have switched off the temperature-dependent thermal pressure when generating the model results. This simplification leads to an overestimation of density and an underestimation of temperature at low pressure levels and high temperatures by less than 5%. The resulting impact on the planetary radius is minimal, typically on the order of 0.1%, which is insignificant since the overall radius of most sub-Neptunes is primarily determined by the H/He envelope and radiative atmosphere, rather than the rock/iron core.
@@
It is important to note that we have neglected the contribution of gravitational energy release due to the contraction of the iron core. This treatment is based on our finding that the radius contraction rate is minimal over Gyr timescales ($\sim$1%), resulting in gravitational energy release contributing at most 10% of the energy flux at the CMB. This energy release is negligible compared to the cooling energy $L_{\rm{int}}$ that drives envelope contraction.
@@
where $T_{\rm{CMB,c}}$, $\alpha_{\rm{CMB,c}}$, $\rho_{\rm{CMB,c}}$ are the temperature, thermal expansivity and density of the iron core at the CMB. The second equality holds when the potential temperature $T_{c}$ is defined at the CMB pressure for each timestep for computational convenience. The choice of this reference pressure does not influence the final result. By tracking $T_{\rm{CMB,c}}$, we can accurately evolve the iron core’s thermal state along the adiabat. The second term in the middle, associated with the pressure change at the CMB, $\dot{P}_{\rm{CMB}}$, is updated during each hydrostatic calculation. Although this pressure change typically evolves over a long timescale and contributes minimally to the energy budget during the thermal contraction phase, it becomes significantly impactful during periods of boil-off. Rapid depressurization due to the removal of envelope mass can substantially influence the interior energy budget, highlighting the necessity of this decompression term.
@@ 2.7 Silicate Mantle
For the solid mantle, the dominant mineral types vary with pressure. Below 25 GPa, the lower mantle consists primarily of Mg-perovskite ($\rm{MgSiO_{3}}$), while the upper solid mantle is composed of olivine ($\rm{Mg_{2}SiO_{4}}$). For simplicity, we divide the solid-phase mantle into two distinct layers to account for the pressure-dependent variation in the Mg:Si ratio, each with a fixed composition. However, it is important to acknowledge that a more complex composition, including a variable Mg:Si ratio and the presence of chemical impurities, could significantly influence interior dynamics, resulting in varied thermal evolution. Addressing these complexities will be an important focus for future studies.
The EOS for Mg-perovskite is based on Keane’s formulation (Keane, 1954), which reduces uncertainties at high pressures compared to the BM3 model:
@@
where $K_{0}$, $K^{\prime}_{0}$, $K_{0}^{\prime\prime}$ and $\rho_{0}$ are the model parameters. For the upper mantle, the EOS follows the BM3 formulation (Eq. 25).
@@
The solidus for the upper mantle (pressures $<$ 10 GPa) is based on experiments on peridotite (Hirschmann, 2000). Since the liquidus data from the same experiments carry significant uncertainty, we assume the liquidus is 200 K above the solidus. At higher pressures, the solidus and liquidus are derived from experimental data on chondrites (Andrault et al., 2011). For pressures exceeding 300 GPa, where experimental data is unavailable, the liquidus is taken from the ab initio melting curve of pure Mg-perovskite (Belonoshko et al., 2005), with the solidus assumed to be 85% of the liquidus temperature. The T-P relationship of the solidus and liquidus are shown in the bottom right panel of 3. We integrate the local melt fraction $\phi$ over each mass shell to determine the total masses of the solid and molten silicates, denoted as $M_{s}$ and $M_{m}$, respectively.
The temperature difference across the interface between the two solid layers (the upper and lower solid mantle) is assumed to be zero, meaning they evolve together as a single thermodynamic entity. We evolve the magma ocean adiabat and solid mantle adiabat separately due to their contrasting rates of heat transport. The temperature jump at the PCI governs the energy transfer rate between the solid and liquid layers. Nevertheless, our results indicate that this temperature jump is relatively small ($\lesssim$ 100K) compared to that at the CMB (exceeding 1000K). Consequently, the mantle bottom, rather than the PCI, acts as the rate-limiting layer for energy transport in the mantle.
@@
where $M_{\rm{Si}}=M_{s}+M_{m}$ represents the total silicate mass. In these equations, the rate of change in the potential temperatures of the magma ocean, $T_{m}$, and solid mantle, $T_{s}$, are evaluated similarly to Eq. 27. $T_{c}$, $T_{m}$ and $T_{s}$ are evolved simultaneously with $s$ and $M_{\rm{env}}$ using the ODE solver, based on Eq. 26, 31 and 32. On the right hand sides, the energy transport rates ($L_{\rm{CMB}}$, $L_{\rm{PCI}}$ and $L_{\rm{MEB}}$) at each boundary layers are included, alongside heat production from silicate solidification $L_{\rm{Si}}=l_{\rm{Si}}\dot{M}_{\rm{s}}$ and radioactive decay $L_{\rm{radio}}$. Similar to the iron core, we find that the gravitational energy release from mantle contraction contributes minimally to heating. An example of the evolutionary trajectory of these luminosities is shown in the top right panel of Figure 3. $L_{\rm{radio}}$ is evaluated with:
@@
The heat transport process in the mantle is dominated by mantle convection. To evaluate the luminosity at each boundary layer, we employ parameterized convection, which models the process using scaling laws derived from fluid dynamics instead of solving the full set of fluid dynamic equations. This approach has been widely used for Earth and other solar system bodies (e.g., Turcotte et al., 1979; McKenzie & Richter, 1981; Nimmo & Stevenson, 2000; Nimmo et al., 2004).
@@
where $k$ is thermal conductivity of the mantle, whose value varies depending on the physical state and the pressure level. $\Delta T_{\rm{CMB}}=T_{\rm{CMB},c}-T_{\rm{CMB},m}$ sets the temperature gradient at the boundary. The thickness of the thermal boundary layer $\delta$ is parameterized using scaling laws. Two regimes are considered here. The mantle convects in the so-called “hard turbulence” regime (Shraiman & Siggia, 1990) in the liquid state ($\phi>\phi_{\rm{crit}}$), because of the low viscosity and large thickness of the magma ocean. For the solid convection ($\phi\leq\phi_{\rm{crit}}$), $\delta$ is evaluated using the scaling law for the local Rayleigh number $Ra$, which approaches the critical Rayleigh number $Ra_{c}$ at the thermal boundary (Nimmo & Stevenson, 2000; Nimmo et al., 2004). These give:
$\delta=\begin{dcases}d/Nu=d/(0.3Ra^{2/7}),&\text{if }\phi>\phi_{\rm{crit}}\\ \left[\frac{Ra_{c}\kappa_{m}\eta}{\rho g\alpha\Delta T_{\rm{CMB}}}\right]^{1/3},&\text{if }\phi\leq\phi_{\rm{crit}}\\ \end{dcases}$ (35)
@@
where $\rho$ is local density, $g$ is local gravity, $\alpha$ is local thermal expansivity, $\kappa_{m}$ is the thermal diffusivity of the mantle, $\eta$ is the dynamic viscosity of the mantle, and $d$ is the thickness of the magma ocean. The critical melt fraction $\phi_{\rm{crit}}$ is chosen to be 0.5, inspired by Costa et al. (2009); Bower et al. (2018). The Nusselt number $Nu$ parameterizes the increased heat transfer due to convection and depends on $Ra$.
@@
The parameters $\eta_{0}$, $E_{a}$, $V_{a0}$ and $T_{0}$ are provided in Table 2.7. The transition from the low-viscosity magma (with $100\ \rm{Pa\ s}$) to the highly viscous solid phase (with $10^{22}\ \rm{Pa\ s}$) is quite abrupt. To capture this behavior, we adopt a hyperbolic switch function (Bower et al., 2018) $w$ with a transition width $\Delta\phi=0.1$ and a critical melt fraction $\phi_{\rm{crit}}=0.5$: